# Open quantum battery in the background of a three-dimensional rotating black hole

### v1

@@ III.1 Dynamics of the average energy
In Fig. 2 the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ of the quantum battery shown in (14) is plotted as a function of the charging time. We fix the charging amplitude as $A\ell=0.1$, the rotating BTZ black hole mass as $M=1$, and the location of the quantum battery as $R=1.01$. Representative examples of different energy-level spacing of the quantum battery $\Delta\ell=0.05$ (Fig. 2 (a), (b)), and $\Delta\ell=0.15$ (Fig. 2 (c), (d)) have been chosen. We also plot the maximum average energy stored in the closed quantum battery driven by a static driving protocol (corresponding to $\lambda\sqrt{\ell}=0$ case in the interacting Hamiltonian given in Eq. (II)) as a reference limit (see black dotted curves), which is given by $\langle E_{\mathrm{B}}(\tau)\rangle_{\text{static}}=\Delta\frac{A^{2}}{\Omega^{2}}$.
When the energy level spacing $\Delta\ell$ of the quantum battery is smaller than the charging amplitude $A\ell$, under Dirichlet boundary conditions with dissipation (Fig. 2(a)), stored in the quantum battery exhibits damped oscillations over time, with amplitude decaying exponentially due to incoherent relaxation and dephasing rates Eq. (III.1). Dissipation decreases the maximum average energy in an open quantum battery compared to a closed one. Higher angular momentum leads to greater maximum average energy. In limited charging time, increasing the black hole’s rotational speed via dissipation can boost the average energy stored in the quantum battery. However, in the infinite time limit, the average energy converges to a constant, which is given by Eq. (18). In addition, higher angular momentum leads to lower asymptotic average energy in the quantum battery. For transparent boundaries, it reduces both the maximum and asymptotic average energy during charging, unlike Dirichlet boundaries. Thus, increasing black hole rotation decreases stored energy under transparent boundary conditions.
@@
If $\Delta\ell>A\ell$, we can see from Fig. 2(c) and Fig. 2(d) that the average energy exhibits a quite different behavior compared with the $\Delta\ell<A\ell$ case shown in the Fig. 2(a) and Fig. 2(b). Notably, the maximum and asymptotic average energy stored in the open quantum battery can exceed that of the closed quantum battery during the charging process. This indicates the quantum battery extracts energy not only from the charger but also from external vacuum fluctuations in rotating BTZ spacetime, similar to the non-rotating BTZ black hole Tian et al. (2025). Additionally, lower angular momentum increases the stored average energy, meaning black hole rotation suppresses energy extraction and weakens charging performance. This applies to both Dirichlet and transparent boundary conditions. Note: whether the open quantum battery outperforms the closed one depends on the black hole’s rotation.
@@
The behavior of the average energy also depends on the boundary condition of the quantum field at the spatial infinity very much. Compared Fig. 2(a) with Fig. 2(b), we can find that the boundary condition may induce absolutely opposite behavior of the average energy when varying the angular momentum for the $\Delta\ell<A\ell$ scenario. In these two boundary condition cases, they have different maximum and asymptotic evolutionary dynamics. However, for the $\Delta\ell>A\ell$ case, the transparent boundary condition and the Dirichlet boundary condition lead to the same behavior of the average energy via the charging time by varying the angular momentum, while quantitatively different. Moreover, the greater the angular momentum $J/M\ell$, the more pronounced the effect of the boundary conditions on the dynamical behavior of the average energy. However, although the average energy stored in the quantum battery during finite-time charging exhibits different evolution for different boundary conditions, in the infinite-time limit (18), the stored energy depends only on the rotation of the BTZ black hole and is independent of the boundary conditions.
@@
In Fig. 3 we plot the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ stored in the quantum battery as a function of the angular momentum. It is found that with the increase of angular momentum, the maximum average energy increases monotonously for the Dirichlet boundary condition case. However, for the transparent boundary case the maximum average energy decreases monotonously when the angular momentum increases. Furthermore, for the extreme case of the rotating BTZ black hole, i.e., the angular momentum approaches to $1$, the average energy, for both the Dirichlet and the transparent boundary cases, is quite sensitive to the variation of the angular momentum. This means that as the black hole approaches the extremal limit, even small changes in the rotational speed can trigger abrupt changes in the amount of maximum average energy stored in the quantum battery.
@@ III.2 Effect of dissipation on charging power
In Fig. 4 we plot the average charing power as a function of the charging time with varying the angular momentum while fixing other parameters. When $\Delta\ell<A\ell$, the average charing power increases with angular momentum under the Dirichlet boundary conditions but decreases under the transparent boundary conditions. This indicates that black hole rotation enhances charging performance for Dirichlet boundaries while weakening it for transparent boundaries. These are shown in Fig. 4 (a) and (b). When $\Delta\ell>A\ell$, we can see from Fig. 4 (c) and (d) that for both the Dirichlet and transparent boundary conditions, the average charing power is weakened by the angular momentum, which is quite different from the $\Delta\ell<A\ell$ case. This means that, in the case $\Delta\ell>A\ell$, the boundary conditions do not affect the degradation of the charging performance of the quantum battery caused by the rotation of the black hole.
@@
In Fig. 5, we show how the maximum charging power $P_{\max}$ is affected by the angular momentum. When $\Delta\ell<A\ell$, it is shown in Fig. 5 (a) that the maximum charging power increases monotonously with the increase of the angular momentum. Near the extreme angular momentum, the maximum charging power becomes highly sensitive to small changes in the black hole’s rotational speed, significantly improving the quantum battery’s charging performance under the Dirichlet boundary condition. However, under the transparent boundary condition, the maximum charging power decreases monotonically with increasing angular momentum and becomes more sensitive near the extreme condition, as shown in Fig. 5 (b). When $\Delta\ell>A\ell$, the maximum charging power decreases monotonously with the increase of the angular momentum for both the Dirichlet and transparent boundary condition cases, as shown in Fig. 5 (c) and (d). However, the maximum charging power for the Dirichlet boundary condition case decrease faster than that for the transparent boundary condition case when the angular momentum is relatively small, while decreases more slowly when the angular momentum is near the extreme case. This further reveals the distinct effects of boundary conditions in the context of a quantum battery whose charging performance is influenced by the rotation of a black hole.
@@ III.3 Energy flow in quantum battery
From the above analysis, we find that whether the open quantum battery outperforms the closed one depends on the ratio of energy level spacing to charging amplitude. To better understand this charging performance physics, we investigate the flow of average energy during the charging process, as shown in Fig. 6. It is found that when $\Delta\ell<A\ell$, curves associated with the quantum battery and the charger are not specular with respect to the $x$-axis, this reflects that a certain amount of energy is dissipated into the reservoir. Similarly, the curves associated with the quantum battery and the charger are also not specular with respect to the $x$-axis when $\Delta\ell>A\ell$. And the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ change associated with the quantum battery is greater than the average energy $\langle E_{\mathrm{C}}(\tau)\rangle$ change associated with the charger, which reflects the fact that the average energy of the quantum battery is not only obtained from the charger but also extracted from the vacuum fluctuations of the quantum fields in rotating BTZ black hole.
Moreover, from the dynamics of the average energy evolution, it is evident that the dissipation plays a dominant role in the overall charging process at this stage, as shown in Fig. 6(c) and (d). It is worth noting that according to the variation of the average energy dynamics curves, the quantum battery would extract less energy from the vacuum fluctuations when the angular momentum is larger compared to when the angular momentum is smaller. This means that the rotation of the BTZ black hole inhibits the quantum battery from extracting energy from the vacuum fluctuations.

### v3

@@ III.1 Dynamics of the average energy
Fig. 2 shows the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$, given by Eq. (13), stored in the quantum battery as a function of the evolution time $\tau/\ell$. We plot the maximum average energy stored in the closed quantum battery driven by a static driving protocol (corresponding to $\lambda\sqrt{\ell}=0$ case in the interacting Hamiltonian $H_{I}$ as a reference limit (see black dotted curves), which is given by $\langle E_{\mathrm{B}}(\tau)\rangle_{\text{static}}=\Delta\frac{A^{2}}{\Omega^{2}}$.
From Fig. 2(a), we find that, when the energy level spacing $\Delta\ell$ of the quantum battery is smaller than the charging amplitude $A\ell$, the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ does not exceed that stored in a closed quantum battery for Dirichlet boundary conditions. Moreover, the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ exhibits damped oscillations over time. This is because the energy transferred from the charger to the quantum battery, due to environmental effects, introduces incoherent relaxation rate $\mathcal{K}^{\prime}$ and dephasing rate $\mathcal{K}$, thereby leading to the system’s dissipative evolution. Specifically, the vacuum in the background spacetime of a nonrotating ($J/Ml=0$) BTZ black hole exhibits maximal dissipation. For rotating BTZ black holes ($J/Ml\neq 0$), the amplitude of the energy oscillations decreases as the angular momentum $J/Ml$ increases. This indicates that, because of the interaction between the quantum battery and its environment, the black hole’s nonzero angular momentum can reduce the incoherent relaxation and dephasing rates in Eq. (III.1), thereby suppressing energy dissipation from the dynamics. When the quantum battery is isolated from its environment, its energy exhibits periodic variations, as described by Eq. (16). However, as shown in Fig. 2(b), for transparent boundary conditions, the rotational effects of the black hole accelerate the energy dissipation from the dynamics. This effect is attributed to the fact that nonzero angular momentum increases rather than decreases the incoherent relaxation rate $\mathcal{K}^{\prime}$ and the dephasing rate $\mathcal{K}$ in Eq. (III.1) for this boundary conditions, resulting in greater dissipation.
@@
For $\Delta\ell>A\ell$, we can see from Fig. 2(c) and Fig. 2(d) that the average energy exhibits a quite different behavior compared with the $\Delta\ell<A\ell$ case shown in the Fig. 2(a) and Fig. 2(b). Notably, the maximum average energy stored in the open quantum battery can exceed that of the closed quantum battery during the charging process. This suggests that the quantum battery can obtain energy not only from the charger but also from vacuum fluctuations via dissipative processes, similar to the static BTZ black hole Tian et al. (2025). Notably, in the static BTZ spacetime (i.e., $J/Ml=0$), the quantum battery attains its maximum stored energy, indicating maximal extraction of energy from vacuum fluctuations for this conditions. However, as the black hole starts rotating (i.e., as its angular momentum $J/Ml$ gradually increases), we observe that the energy stored in the open quantum battery decreases. For $J/Ml<0.9$, the energy stored in the open quantum battery falls below that of the closed quantum battery. These results indicate that, as the BTZ black hole transitions from a non-rotating (static) state to a rotating state, an increase in the rotation parameter suppresses energy extraction and thus reduces the quantum battery’s energy storage capacity. This outcome arises because the black hole’s nonzero angular momentum modifies vacuum fluctuations, thereby modulating the quantum battery’s energy dissipation rate and affecting the battery’s performance. This applies to both Dirichlet and transparent boundary conditions.
@@
The behavior of the average energy also depends on the boundary condition of the quantum field at the spatial infinity very much. Compared Fig. 2(a) with Fig. 2(b), we can find that the boundary condition may induce absolutely opposite behavior of the average energy when varying the angular momentum for the $\Delta\ell<A\ell$ scenario. In these two boundary condition cases, they have different maximum and asymptotic evolutionary dynamics. However, for the $\Delta\ell>A\ell$ case, the transparent boundary condition and the Dirichlet boundary condition lead to the same behavior of the average energy via the charging time by varying the angular momentum $J/Ml$, while quantitatively different. Moreover, the greater the angular momentum $J/M\ell$, the more pronounced the effect of the boundary conditions on the dynamical behavior of the average energy. However, although the average energy stored in the quantum battery during finite-time charging exhibits different evolution for different boundary conditions, in the infinite-time limit in Eq.(17), the stored energy depends only on the rotation of the BTZ black hole and is independent of the boundary conditions.
@@
In Fig. 3 we plot the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ stored in the quantum battery as a function of the angular momentum $J/Ml$ for $\Delta\ell<A\ell$. We observe that the variation in the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ depends strongly on the boundary conditions. Specifically, with the increase of angular momentum $J/Ml$, the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ increases monotonously for the Dirichlet boundary condition case. However, for the transparent boundary case the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ decreases monotonously when the angular momentum $J/Ml$ increases. These two boundary conditions have opposite effects on changes in energy, as in the static black hole scenario Tian et al. (2025). Furthermore, for the extreme case of the rotating BTZ black hole, i.e., the angular momentum $J/Ml$ approaches to $1$, the average energy, for both the Dirichlet and the transparent boundary cases, is quite sensitive to the variation of the angular momentum $J/Ml$. This means that as the black hole approaches the extremal limit, even small changes in the rotational speed can trigger abrupt changes in the amount of maximum average energy stored in the quantum battery. This behavior may result from dramatic variations in the vacuum state near the extremal black hole.
@@ III.2 Effect of dissipation on charging power
In Fig. 4 we plot the average charing power $P_{\mathrm{B}}(\tau)$ as a function of the charging time $\tau/\ell$ with varying the angular momentum $J/M\ell$ while fixing other parameters. In each image, the initial peak of the average power was the largest, and the peak decreased gradually over time. Initially, energy can flow into the battery at the maximum possible rate. However, subsequent oscillations arise because the system is open and interacts with the vacuum field. This interaction induces irreversible dissipation and decoherence. Consequently, vacuum fluctuations act as a noisy environment, causing the system to lose coherence and thereby reducing the charging power. Moreover, we observe in Fig. 4(a) that, when $\Delta\ell<A\ell$, the charging performance of the quantum battery is minimized in the static BTZ black hole background ($J/M\ell$ =0) with the Dirichlet boundary condition. However, in the rotating BTZ black hole background ($J/M\ell\neq 0$), increasing the angular momentum $J/Ml$ enhances the battery’s $P_{B}(\tau)$. This results imply that black hole rotation enhances quantum battery performance for the considered condition. For other conditions (see Figs.4 (b)–(d)), increasing the black hole’s angular momentum $J/Ml$ reduces the $P_{B}(\tau)$ of the quantum battery.
@@
To obtain a comprehensive understanding of the impact of black hole rotation on the charging process of the quantum battery, we plot the maximum charging power $P_{\max}$ as a function of the angular momentum $J/M\ell$ in Fig. 5. As shown in Fig. 5(a), for Dirichlet boundary conditions and when $\Delta\ell<A\ell$, the maximum charging power $P_{\max}$ of the quantum battery increases monotonically with the angular momentum $J/M\ell$. This demonstrates that, for these conditions, optimal charging power is obtained for vacuum fields around rapidly rotating black holes (i.e., black holes with high angular momentum). Notably, unlike Fig. 5(a), Figs. 5(c)-(d) show that the maximum charging power $P_{\max}$ of the quantum battery decreases monotonically as the angular momentum $J/Ml$ increases. The results imply that the quantum battery achieves optimal charging power when coupled to the vacuum field in the background spacetime of a static (non-rotating) BTZ black hole. Moreover, this optimal charging power cannot be increased by increasing the angular momentum $J/M\ell$. Specifically, when the black hole approaches the extremal limit, where $J/M\ell\to 1$, the charging power varies sharply. This may be attributed to drastic changes in the quantum vacuum as the black hole approaches extremality.
@@ III.3 Energy flow in quantum battery
From the above analysis, we find that whether the open quantum battery outperforms the closed one depends on the ratio of energy level spacing $\Delta\ell$ to charging amplitude $A\ell$. To better understand this charging performance physics, we investigate the flow of average energy during the charging process, as shown in Fig. 6. It is found that when $\Delta\ell<A\ell$, curves associated with the quantum battery and the charger are not specular with respect to the $x$-axis, this reflects that a certain amount of energy is dissipated into the reservoir. Similarly, the curves associated with the quantum battery and the charger are also not specular with respect to the $x$-axis when $\Delta\ell>A\ell$. And the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ change associated with the quantum battery is greater than the average energy $\langle E_{\mathrm{C}}(\tau)\rangle$ change associated with the charger, which implies the fact that the average energy of the quantum battery is not only obtained from the charger but also extracted from the vacuum fluctuations of the quantum fields in rotating BTZ black hole, this result is consistent with the conclusion presented above.
In particular, from the dynamics of the average energy evolution, it is evident that the dissipation plays a dominant role in the overall charging process, as shown in Fig. 6(c) and (d). Notably, when the angular momentum $J/Ml=0$, i.e., for a stationary BTZ black hole, the quantum battery extracts maximal energy from vacuum fluctuations via dissipative processes. However, as the angular momentum $J/Ml$ increases, i.e., as the BTZ black hole begins to spin, the energy the quantum battery extracts from vacuum fluctuations gradually decreases. This means that the rotation of the BTZ black hole inhibits the quantum battery from extracting energy from the vacuum fluctuations.